Three point charges are arranged on a line. Charge q3 = +5.00 nC and is at the origin. Charge q2 = -3.27 nC and is at x = +4.47 cm. Charge q1 is at x = +2.49 cm. What is q1 (magnitude and sign) if the net force on q3 is zero? (Give your answer in decimal using nC as unit)
Answer #1

in order to get the null point at q3, electric field at that chargedue to q2 and q1 must be equal to zero.

now

k q1 / r1^2 = k q2 / r2^2

q1 = q2 ( r1/r2) ^2

q1 = 3.27* ( 2.49 / 4.47)^2

q1 = 1.015 nC

=========

Comment in case any doubt, will reply for sure.. Goodluck